OpenStudy (anonymous):

A bucket of tennis balls contains 12 yellow, 5 orange, and 3 pink tennis balls. You randomly choose a ball put it back, and randomly choose another ball. What is the probability that the first ball is yellow and the second ball is orange?

OpenStudy (anonymous):

Show me your attempt =)

OpenStudy (anonymous):

Add up how many Balls you have, 12+5+3 = 20 Balls There are 12 Yellow balls Your probability of a yellow ball is 12/20 = 6/10 = 60% chance that your first ball is yellow. You put the ball back, There are 5 orange Balls, 5/20 = 1/4 = 25% chance your next ball is Orange.

OpenStudy (anonymous):

the second should still be yellow because there are very many yellows

OpenStudy (anonymous):

thanks ^ but i want to learn how to do it by myself do i can know how to do it later on on my own :)

OpenStudy (anonymous):

That's mah girl =)

OpenStudy (anonymous):

so i know there's 12/20 yellow and 5/20 orange

OpenStudy (anonymous):

do i use that technique you told me before or is this different? @SmoothMath

OpenStudy (anonymous):

Totally the same. I have two events, right? And I want to know the probability of both of them happening, right?

OpenStudy (anonymous):

but we want to know what is the probability that the first ball is yellow and then the second one is orange

OpenStudy (anonymous):

aren't they suppose to be seperate?

OpenStudy (anonymous):

Not sure what you mean by separate?

OpenStudy (anonymous):

because we want to know the probability of yellow first. and then next the orange so isn't there should be 2 separate fractions?\

OpenStudy (anonymous):

Riiiight. The first event, pulling a yellow tennis ball. What is the probability of that happening? The second event, we put all the balls back in, and we try to pull an orange tennis ball. What is the probability of that happening?

OpenStudy (anonymous):

12/20 and then 5/20

OpenStudy (anonymous):

Coooool. So the probability of the first thing and THEN the second thing? Whatdoyoudo?

OpenStudy (anonymous):

multiply them together so it will be 60/20?

OpenStudy (anonymous):

Well, when you look at that probability, you see a number greater than 1. That should concern you. Something that is certain to happen has a probability of 1. There's no such thing as a probability higher than that. So what mistake did you make? You did want to do (12/20)*(5/20) That part was correct.

OpenStudy (anonymous):

Perhaps you forgot how to multiply fractions? That happens a lot.

OpenStudy (anonymous):

wait so we should multiply the 20 and 20 also?

OpenStudy (anonymous):

so it would be 60/400 ?

OpenStudy (anonymous):

which is also 3/20 or 0.15

OpenStudy (anonymous):

=D

OpenStudy (anonymous):

yay :D

OpenStudy (anonymous):

Okay, fun time question. What would the probability be of pulling a yellow and then an orange if I DIDN'T put the yellow ball back in after?

OpenStudy (anonymous):

oh um give me a second to work it out

OpenStudy (anonymous):

60/380?

OpenStudy (anonymous):

12/20 * 5/19

OpenStudy (anonymous):

@SmoothMath are you there????

OpenStudy (anonymous):

You nailed it. You're a champion.

OpenStudy (anonymous):

lol thanks! :) i have 2 others, sorry to bother you. but can you help me on those also? i'll post it into a new question. sorry

OpenStudy (anonymous):

Do not apologize. I'm sure you can do these though.

OpenStudy (anonymous):

ok but can you help me if i get stuck cause the next one has so many numbers that i do not know where to start